The price elasticity of demand for ambulatory mental health services appears to be about - 0.8, and the price elasticity for general ambulatory medical services appears to be about -0.3. How much would spending increase for each type of care if copays were cut from $40 to $25? *Hint: Need to find out how much the copay decreased by in percentage and that will give you the percentage change in price to plug into your elasticity formula. (PLEASE SHOW WORK ON HOW YOU GOT THE ANSWER, THANK YOU).
Answer #1

Given, The price elasticity of demand for ambulatory mental health services is -0.8. The price elasticity of demand for generFor general ambulatory medical services, -0.3 Percentage change in quantity demanded / -37.5% Percentage change in quantity dTherefore, new spending would be (0.625P) (1.1125Q) 0.6953125Q. This shows that new spending has decreased, and it decreased

Given, The price elasticity of demand for ambulatory mental health services is -0.8. The price elasticity of demand for general ambulatory medical services is -0.3. If copay were cut from $40 to $25, that is price decreased, the percentage change in price can be computed as, [($25 - $40) $40] -37.5% Using the formula for price elasticity of demand, use the values above to get the change in quantity demanded Price elasticity of demand = (Percentage change in quantity demanded / Percentage change in Prices) For ambulatory mental health services, Percentage change in quantity demanded / -37.5% -0.8 Percentage change in quantity demanded (0.8 * 37.5%) = 30% With the new prices, quantity demanded has increased by 30%.

For general ambulatory medical services, -0.3 Percentage change in quantity demanded / -37.5% Percentage change in quantity demanded = (0.3 *37.5%)= 11.25% With the new prices, quantity demanded has increased by 11.25% Though spending on these services actually decreased, as it can be shown by multiplying the new price with new quantity demanded. For ambulatory mental health services, Taking old prices as P and old quantity demanded as Q, new =0.625P. New quantity demanded would be (1 0.3) * Q = 1.3Q. prices would be (1 -0.375) * P Therefore, spending has decreased, and it decreased by (1 0.8125)* 100 18.75%. new spending would be (0.625P) (1.3Q) 0.8125PQ. This shows that new For general ambulatory medical services, Taking old prices as P and old quantity demanded as Q, new =0.625P. New quantity demanded would be (1 0.1125)* Q prices would be (1 -0.375) * P 1.1125Q. #

Therefore, new spending would be (0.625P) (1.1125Q) 0.6953125Q. This shows that new spending has decreased, and it decreased by (1 - 0.6953125)* 100 = 30.46875%, that is around 30.47%.